An invester is concerned about interest rate risk and decides to lock in a fixed interest rate on January 1, 2011. The notional balance is $21500, the fixed rate is a nominal rate of 4.8% compounded semiannually.
The floating rate is the six-month LIBOR, and the swap is for eighteen months. The LIBOR rates, given on an Actual/Actual basis turn out to be 2.583% for the period beginning January 1, 2011, 4.664% for the period beginning July 1, 2011, and 3.345% for the period beginning January 1, 2012. Determine the amount of the payments (from the invester's standpoint) at the end of each six-month period.

Note: The stated LIBOR rates are computed as simple interest!

(a) The payment at the end of the first six months is $ ________________

(b) The payment at the end of the second six months is $ _______________

(c) The payment at the end of the third six months is $___________________
Answer #1

a]

Payment = notional amount * (fixed rate - floating rate) * 6 / 12 = $21,500 * (4.8% - 2.583%) * 6 / 12 = $238.33

b]

Payment = notional amount * (fixed rate - floating rate) * 6 / 12 = $21,500 * (4.8% - 4.664%) * 6 / 12 = $14.62

c]

Payment = notional amount * (fixed rate - floating rate) * 6 / 12 = $21,500 * (4.8% - 3.345%) * 6 / 12 = $156.41